# Generative Verifiers: Reward Modeling as Next-Token Prediction

### v2

@@ Abstract
Verifiers or reward models are often used to enhance the reasoning performance of large language models (LLMs). A common approach is the Best-of-N method, where N candidate solutions generated by the LLM are ranked by a verifier, and the best one is selected. While LLM-based verifiers are typically trained as discriminative classifiers to score solutions, they do not utilize the text generation capabilities of pretrained LLMs. To overcome this limitation, we instead propose training verifiers using the ubiquitous next-token prediction objective, jointly on verification and solution generation. Compared to standard verifiers, such generative verifiers (GenRM) can benefit from several advantages of LLMs: they integrate seamlessly with instruction tuning, enable chain-of-thought reasoning, and can utilize additional test-time compute via majority voting for better verification. We demonstrate that GenRM outperforms discriminative, DPO verifiers, and LLM-as-a-Judge, resulting in a $16-40$% improvement in the number of problems solved with Best-of-N on algorithmic and math reasoning tasks. Furthermore, we find that training GenRM with synthetic verification rationales is sufficient to pick out subtle errors on math problems. Finally, we demonstrate that generative verifiers scale favorably with model size and inference-time compute.
@@ 1 Introduction
[…]
@@
Figure 2: Example using generative CoT verifier on GSM8K test. LLM-generated solutions often sound convincing even when they are wrong, making verification a challenging task. Here, the solution is incorrect because it has ignored the word ‘each’ in the problem. While the discriminative RM fails to recognize this subtle mistake in the solution, our GenRM-CoT verifier reliably detects the error. This is because GenRM-CoT was trained with next-token prediction on synthetic chain-of-thought rationales, enabling it to explicitly reason about the solution. The full verification output can be found in Table E.12.
@@ 3.3 Chain-of-Thought Verifiers (GenRM-CoT)
Synthetic verification CoT rationales for training. Verifying LLM solutions with human-generated rationales can become increasingly expensive and challenging as LLMs surpass human reasoning abilities. To address this challenge, we explore using synthetically-generated rationales on GSM8K. One naive approach is to simply use the ‘Let’s verify step by step’ prompt given a problem-solution pair, and keep the generated rationales only when they accurately verify the correctness of a solution (Singh et al., 2023; Zelikman et al., 2022). However, such rationales (after filtering based on final yes/no responses) are still often of poor quality, due to 50% accuracy from random guessing.
Instead, we use reference-guided grading (Zheng et al., 2024) to improve the quality of synthetic rationales, which we filter using their verification correctness. Specifically, we provide a reference solution in addition to the problem and solution to verify (see Table A.2), making it easier for an LLM to point out any reasoning error in the provided solution. Here, a reference solution could be any model-generated solution that arrives at the correct final answer. Note that reference-guided grading can only be used during training, as we do not have reference solutions for test problems.
@@ 4 Experiments
  • Algorithmic reasoning. We use two difficult string manipulation tasks, namely Last Letter Concatenation (Wei et al., 2022) and Word Sorting from Big-Bench (Suzgun et al., 2022). We train verifiers on word lists of length {2,3,4}, and evaluate their generalization on length {5,6}.
  • Math reasoning. We train grade-school math verifiers on the GSM8K dataset from Cobbe et al. (2021) that popularized test-time verification. We evaluate these verifiers on the GSM8K test set as well as their easy-to-hard generalization on much harder MATH dataset (Hendrycks et al., 2021), using the same held-out set of 500 MATH problems as Lightman et al. (2023).
@@ 4.1 Generative Verifiers Outperform Standard Verification Approaches
On GSM8K, GenRM-CoT consistently outperforms other methods (Figure 7, middle), even though the synthetic CoT rationales for training may contain errors. Qualitatively, GenRM-CoT is able to detect subtle reasoning errors that are missed by discriminative or direct GenRM verifiers (see Figure 2, 4, and 14).
Easy-to-Hard Generalization. Without any training on MATH, GenRM-CoT results in a 6.4×6.4$\times$6.4 × better sample efficiency than discriminative verifiers as we increase the number of solutions to verify, and surpasses the strong self-consistency baseline (Figure 7, right). While Sun et al. (2024) demonstrate that discriminative verifiers trained on easy MATH problems can generalize to harder MATH problems, GenRM-CoT exhibits a much stronger generalization from grade-school math problems to high-school competition problems in MATH (see Figure 7 for a score breakdown by subject areas and difficulty levels).
@@ 4.2 Synergy Between Generation and Verification
Figure 8: SFT on correct solutions enhances verification, both for GenRM and GenRM-CoT, across all tasks. ‘Verification Only’ corresponds to verifiers trained only on verification data, by setting $\lambda=0$ in (5).
@@ 4.3 Scaling Model Size and Inference-time Compute
Scaling model size. In Figure 11, we show that generative verifiers, especially GenRM-CoT, exhibit better scaling behavior than discriminative RMs, both in terms of reward modeling accuracy and Best-of-N performance. Intuitively, bigger models are more capable of text generation, allowing GenRM-CoT finetuning to better tap into its chain-of-thought reasoning ability for verification. Furthermore, these results demonstrate that larger models generalize better using the same data, which matches what we expect from scaling model parameter counts under the next-token prediction loss.
@@ 5 Related Work
LLM-as-a-Judge. Another line of work that poses verification as next-token prediction simply prompts off-the-shelf LLMs to act as a verifier when provided with a rubric and a template for grading (Zheng et al., 2024; Bai et al., 2022; Kim et al., 2023; Ling et al., 2024) or many-shot ICL examples (Agarwal et al., 2024), but without any specific training for the same. Perhaps unsurprisingly, we find in our experiments that using more powerful LLMs (Gemini 1.0 Pro) as a judge is worse than our trained GenRM using weaker Gemma models (Figure 1, 11), highlighting the necessity of training generative verifiers. Our generative verifiers also exhibit good out-of-distribution generalization, which might be due to better calibrated uncertainty estimates from training (Kapoor et al., 2024). More generally, even the strong proprietary LLMs, such as GPT-4 (Achiam et al., 2023) and Gemini  (Team et al., 2024), fall behind trained RMs on popular leaderboards (Lambert et al., 2024), and this gap is much larger for reasoning.
@@
Unified generation and verification. One of the hallmark properties of GenRM is that the same generative verifier can be co-trained with a generation objective (5): when given a problem, the model is trained to produce a solution, whereas when given a problem and a candidate solution, it is trained to verify this candidate. This is related to DPO (Rafailov et al., 2024) and its application to learning verifiers in reasoning (Hosseini et al., 2024), which aims to unify generation (policy) and verification (reward models) by representing the reward implicitly using the logits of a policy and training the policy with a reward-modeling loss. For reasoning, this type of model tying has been shown to exhibit erroneous extrapolation and degradation in learned representations, which prior work has attempted to address with additional techniques (Pang et al., 2024; Setlur et al., 2024; Pal et al., 2024; Yang et al., 2024). Of these, while Yang et al. (2024) train a reward model with an auxiliary generative SFT loss, note that this loss is applied on a separate head for regularization purposes and is discarded after training; unlike GenRM no text is produced when querying the RM. In addition, compared to DPO, GenRM uses a simpler next-token prediction loss, does not require a reference policy, and obtains significantly better verification performance (Figure 1, 7).
@@ 6 Conclusion & Future Work
In this paper, we have introduced Generative Verifiers (GenRM), which recast verification as next-token prediction. GenRM is more performant than discriminative verifiers, and unlocks the use of chain-of-thought reasoning and inference-time compute for better verification. GenRM also unifies generation and verification into a single LLM, and demonstrates that such a unification benefits both generation and verification. Moreover, we show that synthetic model-generated rationales, which can be error-prone, are sufficient to teach GenRM how to use verification CoT to pick out tricky errors on math reasoning tasks (see Figure 2, 4, 14, and Appendix E).
@@ Acknowledgements
LZ led the project, and ran almost all of the experiments and ablation studies, and wrote and edited the paper. AH was responsible for the discriminative RM baselines and DPO baselines. HB was responsible for the word-sorting task, and helped set up evaluations and DPO. MK advised the project, and provided feedback on writing. AK conceived the project with RA, advised LZ, and provided feedback on paper. RA hosted LZ as a student researcher, proposed several ideas and experiments, implemented reference-guided grading, wrote the initial draft and advised the project.

### v3

@@ Abstract
Verifiers or reward models are often used to enhance the reasoning performance of large language models (LLMs). A common approach is the Best-of-N method, where N candidate solutions generated by the LLM are ranked by a verifier, and the best one is selected. While LLM-based verifiers are typically trained as discriminative classifiers to score solutions, they do not utilize the text generation capabilities of pretrained LLMs. To overcome this limitation, we instead propose training verifiers using the ubiquitous next-token prediction objective, jointly on verification and solution generation. Compared to standard verifiers, such generative verifiers (GenRM) can benefit from several advantages of LLMs: they integrate seamlessly with instruction tuning, enable chain-of-thought reasoning, and can utilize additional test-time compute via majority voting for better verification. We demonstrate that GenRM outperforms discriminative, DPO verifiers, and LLM-as-a-Judge, resulting in large performance gains with Best-of-N, namely $5\%\rightarrow 45.3\%$ on algorithmic tasks and $73\%\rightarrow 93.4\%$ on GSM8K. In easy-to-hard generalization settings, we observe improvements of $28\%\rightarrow 44.6\%$ on MATH, and $37.9\%\rightarrow 53.5\%$ on MMLU abstract algebra. Furthermore, we find that training GenRM with synthetic verification rationales is sufficient to pick out subtle errors on math problems. Finally, we demonstrate that GenRM scales favorably with model size and test-time compute.
@@ 1 Introduction
Figure 1: Generative Verifiers outperform standard verification approaches in terms of Best-of-N on reasoning tasks, with a fixed generator. Here, $\Delta$ represents the improvement in number of problems solved with Best-of-N using GenRM-CoT. GenRM-CoT leverages the generation capabilities of LLMs, enabling a finetuned verifier to utilize chain-of-thought verification to detect subtle reasoning errors. For algorithmic tasks, we report average performance using Gemma-2B on Last Letter Concat (Wei et al., 2022) and BBH Word Sorting (Suzgun et al., 2022). For math reasoning, we train Gemma2-9B verifiers on GSM8K and evaluate their performance on GSM8K test (middle) and easy-to-hard generalization on MATH500 (Lightman et al., 2023). For math tasks, LLM-as-a-Judge utilizes Gemini 1.0 Pro, which we used for synthetic verification rationales for training. For each task, the generated solutions in Best-of-N are the same; the only difference is the verifier. Math tasks use model-generated verification rationales for training GenRM-CoT. Data will be released at: https://sites.google.com/view/generative-reward-models.
@@
Figure 2: Example using generative CoT verifier on GSM8K test. LLM-generated solutions often sound convincing even when they are wrong, making verification a challenging task. Here, the solution is incorrect because it has ignored the word ‘each’ in the problem. While the discriminative RM fails to recognize this subtle mistake in the solution, our GenRM-CoT verifier reliably detects the error. This is because GenRM-CoT was trained with next-token prediction on synthetic chain-of-thought rationales, enabling it to explicitly reason about the solution. Note that GenRM-CoT refers to CoT reasoning in the verification process (the solutions typically also contain CoT, but not for verification). The full verification output can be found in Table E.12.
@@ 2 Preliminaries
LLM-as-a-Judge does not finetune a verifier from a pretrained LLM, but simply prompts the LLM to perform the task of verification or self-critique (Zheng et al., 2024; Bai et al., 2022). LLM-judge sometimes uses reference-guided grading, where the LLM is given a reference solution to compare to.
@@ 3.3 Chain-of-Thought Verifiers (GenRM-CoT)
Synthetic Verification CoT Rationales for Training Verifying LLM solutions with human-generated rationales can become increasingly expensive and challenging as LLMs surpass human reasoning abilities. To address this challenge, we explore using synthetically-generated rationales on GSM8K. One naive approach is to simply use the ‘Let’s verify step by step’ prompt given a problem-solution pair, and keep the generated rationales only when they accurately verify the correctness of a solution (Singh et al., 2023; Zelikman et al., 2022). However, such rationales (after filtering based on final yes/no responses) are still often of poor quality, due to 50% accuracy from random guessing.
To improve the quality of synthetic rationales, we provide a reference solution in addition to the problem and solution to verify (see Table A.2), making it easier for an LLM to point out any reasoning error in the provided solution. This idea is similar to reference-guidance grading  (Zheng et al., 2024). Here, a reference solution could be any model-generated solution that arrives at the correct final answer. After initial data generation, we then filter the synthetic rationales using their verification correctness. Note that we condition on a reference solution only to generate training data, but do not include it during actual finetuning of the verifier, so that there is no train/test mismatch.
@@ 4 Experiments
  • Algorithmic reasoning. We use two difficult string manipulation tasks, namely Last Letter Concatenation (Wei et al., 2022) and Word Sorting from Big-Bench (Suzgun et al., 2022). We train verifiers on word lists of length {2,3,4}, and evaluate their generalization on length {5,6}. Note that this is a case of length generalization for the verification task.
  • Math reasoning. We train grade-school math verifiers on the GSM8K dataset from Cobbe et al. (2021) that popularized test-time verification. We evaluate these verifiers on the GSM8K test set as well as their easy-to-hard generalization on much harder MATH dataset (Hendrycks et al., 2021), using the same held-out set of 500 MATH problems as Lightman et al. (2023). We also evaluated model performance on the mathematical tasks in MMLU (Hendrycks et al., 2020) dataset.
@@ 4.1 Generative Verifiers Outperform Standard Verification Approaches
On GSM8K, GenRM-CoT consistently outperforms other methods (Figure 6, middle), even though the synthetic CoT rationales for training may contain errors. Qualitatively, GenRM-CoT is able to detect subtle reasoning errors that are missed by discriminative or direct GenRM verifiers (see Figure 2, 4, and 15).
MMLU Dataset Base Model (Pass@1) Disc-RM GenRM-CoT Improvement
elementary_mathematics 80.1% 90.6% 91.1% +0.5%
high_school_mathematics 52.2% 74.8% 76.1% +1.3%
college_mathematics 47.6% 53% 56.1% +3.1%
abstract_algebra 37.9% 50% 53.50% +3.5%
Table 1: Performance of different methods on math tasks from the MMLU (Hendrycks et al., 2020) dataset. The evaluation uses an easy-to-hard generalization setting, where the verifier is trained only on grade school math. We highlight the absolute improvement of GenRM-CoT over the Disc-RM baseline. Notably, Gen-CoT demonstrates stronger performance across all tasks, with the improvements being more significant on harder tasks.
Figure 7: Transfer to MMLU College Mathematics (GSM Verifiers), using Best-of-32 evaluation, with solutions generated from Gemini 1.0 Pro. On college-level mathematics, even using a single verification rationale with GenRM-CoT can outperform Discriminative RM. Best-of-32 based on discriminative RM yields a performance of $53.0\%$; as for GenRM-CoT (using 32 majority votes), Best-of-32 gives $56.1\%$.
Easy-to-Hard Generalization. Without any training on MATH, GenRM-CoT results in a 6.4×6.4$\times$6.4 × better sample efficiency than discriminative verifiers as we increase the number of solutions to verify, and surpasses the strong self-consistency baseline (Figure 6, right). While Sun et al. (2024) demonstrate that discriminative verifiers trained on easy MATH problems can generalize to harder MATH problems, GenRM-CoT exhibits a much stronger generalization from grade-school math problems to high-school competition problems in MATH (see Figure 6 for a score breakdown by subject areas and difficulty levels) and college-level math in MMLU (see Table 1).
@@ 4.2 Synergy Between Generation and Verification
Figure 9: SFT on correct solutions enhances verification, both for GenRM and GenRM-CoT, across all tasks. ‘Verification Only’ corresponds to verifiers trained only on verification data, by setting $\lambda=0$ in (5). The y-axis of each figure starts from the pass@1 performance of the base generator for each task.
@@ 4.3 Scaling Model Size and Inference-time Compute
Scaling model size. In Figure 12, we show that generative verifiers, especially GenRM-CoT, perform better than discriminative RMs across model sizes, both in terms of reward modeling accuracy and Best-of-N performance. Intuitively, bigger models are more capable of text generation, allowing GenRM-CoT finetuning to better tap into its chain-of-thought reasoning ability for verification. Furthermore, these results demonstrate that larger models generalize better using the same data, which matches what we expect from scaling model parameter counts under the next-token prediction loss.
@@ 5 Related Work
LLM-as-a-Judge. Another line of work that poses verification as next-token prediction simply prompts off-the-shelf LLMs to act as a verifier when provided with a rubric and a template for grading (Zheng et al., 2024; Bai et al., 2022; Kim et al., 2023; Ling et al., 2024) or many-shot ICL examples (Agarwal et al., 2024), but without any specific training for the same. Perhaps unsurprisingly, we find in our experiments that using more powerful LLMs (Gemini 1.0 Pro) as a judge is worse than our trained GenRM using weaker Gemma models (Figure 1, 12), highlighting the necessity of training generative verifiers. Our generative verifiers also exhibit good out-of-distribution generalization, which might be due to better calibrated uncertainty estimates from training (Kapoor et al., 2024). More generally, even the strong proprietary LLMs, such as GPT-4 (Achiam et al., 2023) and Gemini  (Team et al., 2024), fall behind trained RMs on popular leaderboards (Lambert et al., 2024), and this gap is much larger for reasoning.
@@
Unified generation and verification. One of the hallmark properties of GenRM is that the same generative verifier can be co-trained with a generation objective (5): when given a problem, the model is trained to produce a solution, whereas when given a problem and a candidate solution, it is trained to verify this candidate. This is related to DPO (Rafailov et al., 2024) and its application to learning verifiers in reasoning (Hosseini et al., 2024), which aims to unify generation (policy) and verification (reward models) by representing the reward implicitly using the logits of a policy and training the policy with a reward-modeling loss. For reasoning, this type of model tying has been shown to exhibit erroneous extrapolation and degradation in learned representations, which prior work has attempted to address with additional techniques (Pang et al., 2024; Setlur et al., 2024; Pal et al., 2024; Yang et al., 2024). Of these, while Yang et al. (2024) train a reward model with an auxiliary generative SFT loss, note that this loss is applied on a separate head for regularization purposes and is discarded after training; unlike GenRM no text is produced when querying the RM. In addition, compared to DPO, GenRM uses a simpler next-token prediction loss, does not require a reference policy, and obtains significantly better verification performance (Figure 1, 6).
@@ 6 Conclusion & Future Work
In this paper, we have introduced Generative Verifiers (GenRM), which recast verification as next-token prediction. GenRM is more performant than discriminative verifiers, and unlocks the use of chain-of-thought reasoning and inference-time compute for better verification. GenRM also unifies generation and verification into a single LLM, and demonstrates that such a unification benefits both generation and verification. Moreover, we show that synthetic model-generated rationales, which can be error-prone, are sufficient to teach GenRM how to use verification CoT to pick out tricky errors on math reasoning tasks (see Figure 2, 4, 15, and Appendix E).
@@ Acknowledgements
LZ led the project, and ran almost all of the experiments and ablation studies, and wrote and edited the paper. AH was responsible for the discriminative RM baselines and DPO baselines. HB was responsible for the word-sorting task, and helped set up evaluations and DPO. MK advised the project, and provided feedback on writing. AK conceived the project with RA, advised LZ, provided feedback on paper, and helped run additional experiments during rebuttal. RA hosted LZ as a student researcher, proposed several ideas and experiments, implemented reference-guided grading, ran experiments on additional datasets during rebuttal, wrote the initial draft and advised the project.